# Hummus: A Dataset of Humorous Multimodal Metaphor Use

### v1

@@ 1 Introduction
Based on the annotations, we created 6 tasks to test MLLMs’ ability to identify and understand humorous multimodal metaphor use. We tested four open-source and two closed source MLLMs: LLaVA-NeXT-110B/8B, Qwen2-VL-72B/7B-Instruct, GPT-4o, and GPT-4 Turbo, finding that the models struggle with distinguishing metaphorical and non-metaphorical items, as well as understanding the humorous multimodal metaphor use involved. Our ablation study and error analysis reveal that the models’ struggles are likely to be caused by difficulties in integrating visual and textual information.
@@ 2 Background
CMT posits the existence of conceptual metaphors, mappings in the human mind between different conceptual domains Lakoff and Johnson (1980b). Metaphor as a figure of speech is manifestation of conceptual metaphors in language use. Conceptual metaphors also give rise to visual metaphors (e.g., representing an idea as a light bulb) and multimodal metaphors, which involve more than one mode of communication.
@@ 3 Related Works
[…]
@@ 4 Dataset Creation
Hummus is built upon the image-caption pairs in the CapCon corpus Hessel et al. (2023). Each week, The New Yorker publishes a captionless cartoon and receives caption submissions from its readers. The CapCon corpus includes 2578 funny caption submissions for 679 high-quality cartoons of The New Yorker. We randomly sampled 251 cartoons from the corpus, for which a total of 1000 image-caption pairs are available. We designed an annotation scheme informed by theories of humor understanding and metaphor use, and a volunteer linguist with expertise in metaphor annotated the image-caption pairs. Our annotation scheme can be summarized into two stages: humorous metaphor identification and detailed metaphor annotation, which we explain below.
@@ 4.2 Detailed metaphor annotation
5 Data Analysis
@@
Figure 3: Example of (a) unidirectional and (b) bidirectional metaphorical mappings between animals and humans, and example of (c) metaphorically used idiom in our dataset.
@@
Around 390 conceptual metaphors are identified in the dataset. The most frequently occurring ones are animals are humans (23% metaphor samples) and humans are animals (8%). In addition to comparison with animals, the dataset also includes mappings from humans as a source domain to other target domains such as object, god, and aliens. Personification thus appears to be the most frequently used type of metaphor in these humorous image-caption pairs (51%).
@@ 6 Model Evaluation
Task Instructions Evaluation
[…]
The Naming task and the ImageLabel task employ LaBSE Feng et al. (2022) as the evaluator. Our choice is based on a pilot Naming test that involves the first 100 items in our test set and two models: GPT-4o and GPT-4 Turbo. We calculate cosine similarity scores between the model outputs and ground truth using a variety of SBERT models. We choose LaBSE as its predictions are the closest to human judgement.
@@
Table 2: Classification, Naming, and Explanation benchmarks. Standard deviations are in parenthesis. GPT-4 Turbo has a success rate less than 100% in the Naming task (96%).
@@ 6.1 Benchmark results
The random baseline for the Classification task is calculated by randomly choosing between a “Yes” and a “No” answer for each item, and averaging the results of 100 iterations. For ImageBbox, IoU scores above 0.5 are considered a match between model prediction and ground truth. For Naming, ImageLabel, and Explanation, we decide thresholds for good answers by manually examining model answers in various score ranges. The threshold for Naming and ImageLabel is set at 0.6. For Explanation, model answers reaching a ROUGE-1 score of 0.35 and a ROUGE-2 score of 0.087 are considered acceptable.
@@ 6.2 Prompt engineering
Table 4: Mean F1 scores using 3 different prompts for the Classification task. Qwen2 stands for Qwen2-VL-Instruct. Success rate is always 100%.
@@ 6.3 Ablation study
As shown in Figure 4, there is not much difference between the models’ performance in the multimodal and monomodal experiments, except for GPT-4o in the Classification task. The comparison indicates that the multimodal input data is not adequately utilized by the models.
@@ 6.4 Error Analysis
Figure 5: Ground truth versus LLaVA-NeXT-110B predictions for the given image-caption pair. Explanations are shortened. For CaptionHL, the model outputs the entire caption as answer.
@@ 7 Conclusion
We hope this study will encourage more research on MLLMs’ capabilities to process humorous multimodal metaphors. Our dataset is annotated by a single expert. While this ensures efficient, high-quality, and systematic labelling, it may also introduce a certain level of subjectivity to the dataset. We envisage future work exploring alternative interpretations of the image-caption pairs and how MLLMs align with such individuality.
@@ 8 Ethical Considerations
We access the GPT-4 models through OpenAI API. The open-source models (LLaVA-NeXT-8/110B and Qwen2-VL-7/72B) as well as the CapCon corpus are accessed through HuggingFace. The CapCon corpus has CC-BY-4.0 license. Our dataset will be freely accessible on GitHub.
@@
Our dataset includes jokes that could be considered offensive, and certain jokes may be inappropriate for a younger audience. These data remain in our dataset as they can be valuable for future research.

### v2

@@ 1 Introduction
Based on the annotations, we created 6 tasks to test MLLMs’ ability to identify and understand humorous multimodal metaphor use: Classification, Naming (of conceptual metaphors), ImageBbox, ImageLabel, CaptionHL, and Explanation. We tested four open-source and two closed source MLLMs: LLaVA-NeXT-110B/8B, Qwen2-VL-72B/7B-Instruct, GPT-4o, and GPT-4 Turbo, finding that the models struggle with distinguishing metaphorical and non-metaphorical items, as well as understanding the humorous multimodal metaphor use involved. Our ablation study and error analysis reveal that the models’ struggles are likely to be caused by difficulties in integrating visual and textual information.
@@ 2 Background
The Conceptual Metaphor Theory posits the existence of conceptual metaphors, mappings in the human mind between different conceptual domains Lakoff and Johnson (1980b). Metaphor as a figure of speech is manifestation of conceptual metaphors in language use. Conceptual metaphors also give rise to visual metaphors (e.g., representing an idea as a light bulb) and multimodal metaphors, which involve more than one mode of communication.
@@ 3 Related Works
Nevertheless, recent years have seen an increased interest in metaphor understanding. More paraphrase datasets are released Joseph et al. (2023); Tong et al. (2024), following earlier works that treat metaphor understanding as a paraphrasing task Bizzoni and Lappin (2018); Shutova (2010). Other studies frame the issue as an inference task, providing entailed and non-entailed Com s , a et al. (2022); Stowe et al. (2022) or contradictory statements Chakrabarty et al. (2022); Liu et al. (2022b) for metaphorical sentences. A framework for intentions behind metaphor use has also been proposed Michelli et al. (2024).
@@ 4 Dataset Creation
Hummus is built upon the image-caption pairs in the CapCon corpus Hessel et al. (2023). Each week, The New Yorker publishes a captionless cartoon and receives caption submissions from readers. The CapCon corpus includes 2578 funny caption submissions for 679 cartoons of The New Yorker. We randomly sampled 251 cartoons from the corpus, for which a total of 1000 image-caption pairs are available. Each image-caption pair is a unique instance in terms of its humor and possible metaphor use. We designed an annotation scheme informed by theories of humor understanding and metaphor use, and manually annotated the image-caption pairs on Labelbox. Our annotation scheme can be summarized into two stages: (1) humorous metaphor identification, which takes less than a minute per item, and (2) detailed metaphor annotation, which takes $\sim$15 minutes per item.
@@ 4.2 Detailed metaphor annotation
4.3 Inter-annotator agreement
A second expert annotated 300 image-caption pairs randomly sampled from the 1000. Like the first annotator, the second annotator is a linguist with knowledge of CMT and experience in linguistic metaphor annotation.
To ensure timely completion of the task, we left out explanation and figurative device annotation and asked the second expert to (1) classify the items in terms of humorous metaphor use (Yes/No/WIDLII/Discard), (2) identify the conceptual metaphors used in the Yes/WIDLII items, and (3) annotate their image (bounding box) and caption. To facilitate comparison with model performance, we use the same metrics for inter-annotator agreement calculation and model evaluation: multicalss F1 score for humorous metaphor identification, cosine similarity for conceptual metaphor identification, Intersection over Union (IoU) for image annotation, and Jaccard index for caption annotation.
The co-annotation process started with us introducing the second annotator to our annotation scheme. We provided them with the written annotation guidelines and verbally walked them through every step of the scheme (excluding explanation and figurative device annotation), exemplifying both metaphorical and non-metaphorical instances. After the first meeting, the second annotator was assigned the first 50 items to complete independently. Upon finishing the first 50 items, the two annotators inspected their disagreements together and discussed necessary amendments to the original annotation guidelines. After the second meeting, we updated the guidelines and let the second annotator finish the remaining 250 items independently. The annotation was completed in 2 months. With sufficient familiarity with the procedure, annotation for a metaphorical item can be completed in 7 minutes. The guidelines are provided in Appendix A.
The final agreement scores were calculated without further discussion of disagreements: 0.73 average F1 score for humorous metaphor identification, with 0.75 for metaphorical (Yes/WIDLII) items and 0.71 for non-metaphorical (No) items; 0.63 mean similarity score for conceptual metaphor identification ($SD=0.20$); 0.73 mean IoU score for image annotation ($SD=0.32$); 0.65 mean Jaccard index for caption annotation ($SD=0.41$).
Figure 3: Conceptual metaphors identified in our dataset. Font size corresponds to frequency of occurrence.
@@ 4.3 Inter-annotator agreement
Figure 4: Example of (a) unidirectional and (b) bidirectional metaphorical mappings between animals and humans, and example of (c) metaphorically used idiom (go through the roof) in our dataset.
5 Data Analysis
@@
Around 390 conceptual metaphors are identified in the dataset. As shown in Figure 3, The most frequently occurring ones are animals are humans (23% metaphor samples) and humans are animals (8%). In addition to comparison with animals, the dataset also includes mappings from humans as a source domain to other target domains such as object, god, and aliens. Personification thus appears to be the most frequently used type of metaphor in these humorous image-caption pairs (51%).
Task Instructions Evaluation
[…]
@@ 6 Model Evaluation
Table 2: Classification, Naming, and Explanation benchmarks. Standard deviations are in parenthesis. Success rate is 100% except GPT-4 Turbo in the Naming task (96%).
@@
The Naming task and the ImageLabel task employ LaBSE Feng et al. (2022) as the evaluator. Our choice is based on a pilot Naming test that involves the first 100 items in our test set and two models: GPT-4o and GPT-4 Turbo. We calculate cosine similarity scores between the model outputs and ground truth using a variety of SBERT models. We choose LaBSE as its predictions align the most closely to human judgement of good and bad answers.
@@ 6.1 Benchmark results
The random baseline for the Classification task is calculated by randomly choosing between a “Yes” and a “No” answer for each item, and averaging the results of 100 iterations. For ImageBbox, we follow standard practice in object detection Everingham et al. (2010); Lin et al. (2014) and consider IoU scores of 0.5 or higher as indicating sufficient overlap between model prediction and ground truth. For Naming, ImageLabel, and Explanation, we decide thresholds for good answers by manually examining model answers in various score ranges. The threshold for Naming and ImageLabel is set at 0.6. For Explanation, model answers reaching a ROUGE-1 score of 0.35 and a ROUGE-2 score of 0.087 are considered acceptable. Model outputs of different score ranges in these three tasks are presented in Appendix B.
@@ 6.2 Prompt engineering
Table 4: Mean F1 scores using 3 different prompts for the Classification task. Qwen2-VL stands for Qwen2-VL-Instruct. Success rate is always 100%.
@@
Model Classification Naming ImageBbox Caption
Pos Neg Avg
GPT-4 Omni 0.70 0.36 0.53 0.53 (0.23) 0.25 (0.20) 0.50 (0.39)
Turbo 0.73 0.22 0.48 0.50 (0.23) 0.21 (0.20) 0.33 (0.33)
LLaVA-NeXT 110B 0.73 0.00 0.37 0.49 (0.19) 0.17 (0.17) 0.62 (0.43)
8B 0.74 0.01 0.37 0.45 (0.17) 0.15 (0.16) 0.30 (0.31)
Qwen2-VL 72B 0.73 0.00 0.37 0.47 (0.19) 0.17 (0.19) 0.39 (0.39)
7B 0.73 0.00 0.36 0.51 (0.20) 0.13 (0.15) 0.38 (0.36)
Human 0.75 0.71 0.73 0.63 (0.20) 0.73 (0.32) 0.65 (0.41)
Table 5: Model performance compared with human performance on Classification, Naming, ImageBbox (IoU), and Caption (Jaccard index) tasks. Prompts for the models are similar to the annotation guidelines. Some of the tests result in success rate less than 100%: $\geq$90% in Naming; $\geq$90% for the Qwen2-VL models and $\geq$80% for the other models in ImageBbox; $\geq$90% for GPT-4 Turbo and Qwen2-VL-72B, 79% for GPT-4o, and $\sim$60% for the other models in CaptionHL.
6.3 Comparison with human performance
Is the models’ compromised performance really due to the their lack of capacity to process humorous multimodal metaphor use, or is it due to the subjectivity of the tasks? To answer this question, we design a new prompt that is similar to the annotation guidelines we provided to the second expert annotator—the models are instructed to complete Classification, Naming, ImageBbox, and CaptionHL in one go, following the same steps as the expert annotator. We also ask the models to adopt a persona that is representative of the human experts: a linguist who is familiar with CMT and has experience in manual metaphor annotation. The prompt is provided in Appendix B. It is tested on the smallest model, Qwen2-VL-7B, before being used for the full experiment.
As shown in Table 5, even when following similar guidelines, none of the models reach the level of agreement between the two expert annotators. Model performance is closer to human performance in the Naming task, as compared to the other tasks. In CaptionHL, LLaVA-NeXT-110B reaches a Jaccard index score (0.62) that is close to human (0.65), but with a low success rate (62%): for $\sim$40% of the instances the model identifies as metaphorical, it fails to provide a caption annotation. In Classification and ImageBbox, the human expert surpasses the models by a large margin.
Also note that this new approach does not help with the Classification task despite its more detailed instructions about what is metaphor and how to identify humorous metaphor use: most of the models appear to be better at excluding negative cases in the benchmark task (Table 2), although the performance is still much lower than the human expert. Nonetheless, the new approach benefits the Naming task and yields higher Jaccard index scores in the CaptionHL task for GPT-4o, LLaVA-NeXT-110B, and Qwen2-VL-7B, albeit with compromised success rates for the two larger models. Overall, the new approach of following the manual annotation guidelines is better suited for detailed annotation of humorous metaphor use (Naming, ImageBbox, and CaptionHL). As for the Classification task, the models’ inability to identify negative cases seems to go beyond what can be achieved by prompt engineering.
@@ 6.4 Ablation study
As shown in Figure 5, there is not much difference between the models’ performance in the multimodal and monomodal experiments, except for GPT-4o in the Classification task. The comparison indicates that the multimodal input data is not adequately utilized by the models. Our error analysis provides more support for this.
@@ 6.5 Error Analysis
Figure 6: Ground truth (left and in the caption) versus LLaVA-NeXT-110B predictions (right) for the given image-caption pair. Explanations are shortened. For CaptionHL, the model outputs the entire caption as answer.
@@ 7 Conclusion
We hope this study will encourage more research on MLLMs’ capabilities to process humorous multimodal metaphors. While we have experimented with promoting model performance through prompt engineering, other possibilities such as few-shot prompting and fine-tuning are beyond the scope of the current study.
@@ 8 Ethical Considerations
We access the GPT-4 models through OpenAI API. The open-source models (LLaVA-NeXT-8/110B and Qwen2-VL-7/72B) as well as the CapCon corpus are accessed through HuggingFace. The CapCon corpus has CC-BY-4.0 license. Our dataset is freely accessible on GitHub.
@@
Our dataset includes jokes that could be considered offensive, and certain jokes may be inappropriate for a younger audience. These data remain in our dataset as they are a nonremovable part of multimodal humor and can be valuable for future research.
Appendix A Annotation Guidelines
(a) Non-metaphorical.
(b) Metaphorical.
Figure 8: Image-caption pair that (a) does not involve and (b) involves humorous metaphor use.
(a) Correct.
(b) Too small.
(c) Too big.
Figure 9: Good and bad examples for bounding box annotation.
A.1 Objective
You will annotate New Yorker cartoons and their captions in terms of humorous metaphor use.
A.2 Labelling Steps
  1. 1.
    Look at the cartoon image alone (without referring to the caption). Find any incongruities in the image. You can write these down in the incongruity field, but this is optional.
  2. 2.
    Copy and paste the caption to text_hl.
  3. 3.
    How does the caption resolve the incongruities in the image? If you can answer this question, then you understand the humor of the image-and-caption combination. You can write down your understanding in the resolution field, but this is optional.
    1. (a)
      If you understand the humor, continue with the annotation;
    2. (b)
      Otherwise, select “Discard” for the is_met field, assign “1” to complete and submit your annotation for this item.
  4. 4.
    Based on your understanding of the humor, consider whether metaphor use (as detailed in “What counts as metaphor”) is involved in creating that humorous effect. Any kind of involvement counts. In essence, for each metaphor use you find in the image-and-caption combination, consider whether the humor will still be there when the metaphor use is removed.
    1. (a)
      If the answer is an absolute yes, assign “Yes” to is_met and continue with the annotation;
    2. (b)
      If the answer is an absolute no, assign “No” to is_met, “1” to complete and submit your annotation;
    3. (c)
      If you hesitate, assign “WIDLII” to is_met and continue with the annotation.
  5. 5.
    Specify the conceptual metaphor(s) in “X is Y” format. “X” should be the target domain; “Y” should be the source domain. Separate multiple conceptual metaphors with “; ” (a semicolon followed by a whitespace).
    1. (a)
      To decide how general/specific the conceptual metaphor should be, try to relate the instantiation of the metaphor in the image-and-caption combination to other instantiations of the same metaphor (in any modality), and then find a term that can cover them all.
    2. (b)
      You can draw inspiration from this (non-exhaustive) list of conceptual metaphors.
    3. (c)
      Finishing the image_hl and text_hl annotation first (Step 6) could be helpful if you had difficulty pinpointing the two domains.
  6. 6.
    Annotate the metaphor-related image areas (image_hl) and text fragments (text_hl):
    1. (a)
      To annotate metaphor-related image areas, try to assign each object/element of the image to either the target or the source domain of the conceptual metaphors you have annotated. Typically, the majority of the image should belong to the same domain (e.g., the target domain); that one object/element that belongs to the other domain (e.g., the source domain) is the metaphor-related image area. Draw a image_hl bounding box for each such “odd” object/element (e.g., if the minority domain is CAT and there are 3 cats in the image, draw 3 bounding boxes, one box for each cat);
    2. (b)
      Do the same for the caption. In the text_hl field, where you already have the caption copied and pasted, surround each word/phrase/clause that belongs to the “odd” domain with a pair of <i></i> tag. If the entire caption really belongs to a single domain, surround the entire caption in a pair of <i></i> tag. Note that image_hl and text_hl do not have to belong to the same domain.
A.3 Mandatory fields
  1. 1.
    Every item should have a is_met value.
  2. 2.
    If is_met value is “Yes” or “WIDLII”, the following fields should also be filled: image_hl, text_hl (with <i>highlighted text</i>), and conceptual_metaphor.
A.4 What counts as metaphor
Following the Conceptual Metaphor Theory, we define metaphors as conceptual mappings between two difference domains. Therefore:
  1. 1.
    Personification and zoomorphism are considered metaphor use. Personification is metaphor use with HUMAN as the source domain; zoomorphism is metaphor use with ANIMAL as the source domain.
  2. 2.
    Metonymy within the same conceptual domain does not count as metaphor use.
  3. 3.
    Idioms are not metaphors, unless the underlying cross-domain mapping is strictly required to make sense of the humor.
  4. 4.
    Puns can indicate metaphor use. Whether they are metaphorical depends on whether the two meanings can be attributed to some sort of cross-domain mapping.
A.5 Non-metaphorical examples
Example 1: Figure 8(a). Incongruity: Giant beach ball. Resolution: Whoever that was confident to catch the ball is now under it. Why not metaphorical: It’s literally a giant ball; someone’s literally under it in the pool.
Example 2 and 3 are omitted due to space limit.
A.6 Metaphorical examples
Example 1: Figure 8(b). Incongruity: Bear in a restaurant. Resolution: The bear is a waiter. Why metaphorical: Restaurant waiters are human beings. Conceptual metaphor: animals are human workers. Image annotation: See Figure 9(a). Caption annotation: “Thanks for checking on the <i>salmon</i>. How’s the lamb?”.
Example 2 and 3 are omitted due to space limit.
A.7 Bounding box examples
The bounding box should fit tightly around the object of interest—for example, the bear in Figure 9(a). Make sure the entire object is included; Figure 9(b) is therefore incorrect. But also avoid including excessive empty space around the object, such as Figure 9(c).
Appendix B Additional prompts and model outputs
Model Output vs. Ground Truth Score
Dogs are people 0.83
vs. Animals are humans
Temperature is music 0.73
vs. Temperature is pitch
Thoughts are bathwater 0.63
vs. Psychotherapy is a bath
Parenting is egg incubation 0.60
vs. Human baby is an egg
Humor is a drug 0.55
vs. Preaching is a joke
Faith is a wedding 0.45
vs. Alcohol is god
Work is a tool 0.35
vs. Psychotherapy is a bath
Slaying a dragon is a task 0.26
vs. Modern man is knight
Ending a relationship is falling off a cliff vs. Social media is physical world 0.17
Table 6: Sample model outputs in the Naming benchmark task and their cosine similarity scores compared with ground truth.
Model Output Ground Truth Score
Gun Pistol 0.86
Teddy-bear Toy bear 0.75
Meteor Asteroid 0.69
Tree stump Trunk 0.62
Ball Egg 0.57
Hat Cellphone 0.47
Alligator Customer 0.40
Laptop Asteroid 0.32
Beer Human-shaped hole 0.21
Broom Alien spaceship 0.10
Table 7: Sample model outputs in the ImageLabel benchmark task and their cosine similarity scores compared with ground truth.
Figure 10: Sample model outputs in the Explanation benchmark task and their ROUGE-1/ROUGE-2 scores as compared to ground truth.
Figure 11: Prompt for comparing model and human performance.
Table 6, Table 7, and Figure 10 provide example model outputs in the ImageLabel, Naming, and Explanantion benchmark tasks respectively.
The prompt used in Section 6.3 to compare model and human performance is provided in Figure 11.